How many moles of O atoms are present in 88 g of CO2 ?

Grade:11

5 Answers

Ruchi Shaw
9 years ago
Moles = mass / molar mass

Molar mass of CO2 is 44 and given mass of CO2 is 88 gm,
So no. of moles come out to be 2.

Thanks & Regards
Ruchi Shaw
askIITians faculty
Vimlendra Singh
9 years ago
i think 4 moles of O
T C YASHWANTH
9 years ago
total mass co2 is 44 mass of given co2 is 88 in such a way nom of moles we get is 2.
Tapan bapat
9 years ago
CO2 - ATOMIC MASS - 12 + 2 x 16 = 44 . HENCE UNDER - 1 MOLE - I.E UNDER 44 gms - - no of oxygen items are - 02 nos. HENCE - UNDER 44 GMS --- NO OF MOLES OF O ITEM - 2 ( I.E O2) THEREFORE - UNDER 88 GMS - 2 x 2 i.e. 4 moles . IS THE SAID JUSTIFICATION OK.
Aman Kumar Aanand
2 years ago
Given mass of CO2 = 88g
Atomic mass of CO2= 12+16×2= 44g
So no of moles of CO2 in 44g of atomic  mass= 88/44= 2 moles
 
Then no of moles of oxygen in 44g of CO2 = given mass of O2 in CO2/atomic mass of O2
                                                                          16×2/16= 2 moles of O2
And hence, no. Of moles of  02 in 88g of CO2=( 2×88/44) moles= 4 moles of oxygen